OpenStudy (anonymous):

PLEASE HELP write a function that converts h hours to m min. use the function to calulate the number of minutes in a week (168)

OpenStudy (anonymous):

1hours = 60 mins f(h)=h*60 h=hours?? f(168)=168*60=10080